# Measurement-induced transitions for interacting fermions

### v1

@@ VII.1 Summary of main results
In the absence of interaction (and for measurements preserving Gaussianity of the theory), this formalism allows one to identify the symmetry class of the problem in terms of the symmetry of the Lagrangian, Sec. II.6. For complex fermions, this places the problem in the symmetry class AIII or BDI, depending on the presence of additional particle-hole symmetry of Lagrangian, in agreement with earlier results. We have derived the replicated Keldysh non-linear sigma model (NLSM) for both symmetry classes, Sec. III.1. The model incorporates boundary conditions (Fig. 1) specifically designed to produce generating functions for charge cumulants and Rényi entropies directly in the language of the sigma model (Sec. III.2).
@@
We have corroborated these analytical results with a numerical study of an interacting 1D model (Sec. VI). Since exact diagonalization proved insufficient in view of system-size limitations, we have employed the time-dependent variational principle for matrix product states (MPS-TDVP), Sec. VI.1) enabling us to investigate significantly larger systems. The numerically observed behavior of both the entanglement entropy and charge cumulants (Figs. 9 and 10) aligns very well with our analytical predictions. In particular, our MPS-TDVP numerics demonstrate the separation between information and charge, characteristic of the phase with “delocalized” information and charge fluctuations. Furthermore, the behavior of the corresponding behavior of entanglement and charge cumulant is consistent with the predicted area-law and logarithmic-law scaling, respectively. Our estimated values for critical $\gamma$ for both transitions show that two transitions are located quite close to each other. In addition, we have conducted time-dependent Hartree-Fock simulations (Sec. VI.2 and Fig. 11) as a complementary method to estimate the transition point between the area-law phase and the logarithmic phase for charge fluctuations. While the numerical data suggest that the entanglement transition takes place at a somewhat larger value of $\gamma$ than the charge-fluctuation transition (as in the phase diagram in Fig. 8), the accuracy is not sufficient to deduce this with certainty.
@@ VII.2 Comparison with related works
In a number of works [88, 89, 90, 91, 93, 94, 95], the entanglement in 1D models of interacting bosons subjected to generalized measurements was studied numerically. These works reported evidence of a transition from area-law to volume-law scaling of the entanglement entropy. While some of these models were different from our interacting-fermion model, and the measurement protocols were also different, one can expect that the physics that we address has a large degree of universality. In this sense, the conclusions of the above works on the area-law-to-volume-law entanglement transition are in agreement with our results. The scaling of particle-number fluctuations was not considered in these past papers.
@@ VII.3 Outlook
  • (iii)
    Another important direction is to extend the theory to study dynamics (i.e., correlation functions with different time arguments. This will allow, in particular, to directly explore the dynamics of purification and charge sharpening.
@@ Acknowledgements.
We are very grateful to Elmer Doggen, Adam Nahum, Pavel Ostrovsky, Marcin Szyniszewski, and Romain Vasseur for insightful discussions. Note added: While preparing the manuscript for publication, we became aware of related work  [^2] Haoyu Guo, Matthew S. Foster, Chao-Ming Jian, and Andreas W. W. Ludwig, to appear on monitored interacting fermions. We thank Matthew Foster for informing us about this work.
@@ Appendix C Numerical comparison of weak-localization corrections in AIII and BDI classes
In Fig. 12, we plot the behavior of the running coupling constant for both symmetry classes (left panel: AIII, right panel: BDI), as obtained from the density correlation function. The dependence of $g(q)$ at large momenta corresponds to the ballistic behavior: $C(q)\propto\operatorname{const}\Rightarrow g(q)\propto 1/q$; the dashed line marks a universal interpolating function between ballistic ($q\ell_{0}\gtrsim 1$) and diffusion ($q\ell_{0}\ll 1$) regions. Evident bending down of each curve is a manifestation of localization. Clearly, the qualitative behavior of the curves in the two panels of Fig. 12 is identical. At the same, a closer inspection reveals that the localization effects are stronger in the BDI model than in the AIII model, in consistency with analytical predictions
@@ Appendix E Calculation of the Hartree and Fock diagrams for the interaction Lagrangian
As a consequence, at $z\to 0$, the $\tau$-integral in Eq. (217) logarithmically diverges for $d=1$ and saturates at a constant of order unity for $d>1$, yielding Eq. (104) of the main text. The behavior of $Y_{\text{HF}}$ for $d=1,2,3$ is plotted in Fig. 15.
@@ I.1 Time-dependent variational principle
For each bond expansion step, we allow for $k=0.3\chi$ extra singular vectors, expanding a bond of dimension $\chi$. This is somewhat larger than $k=0.1\chi$ discussed in Ref. [135] in the context of DMRG. We found the additional SVD necessary and kept the suggested oversampling parameter of $p\leq 10$. If a maximum bond dimension $\chi_{\rm max}$ is imposed, we still allow for an expansion of any bond beyond $\chi_{\rm max}$ within these expansion parameters at every time step. If the dimension of the expanded bond exceeds $\chi_{\rm max}$ it is truncated back to $\chi_{\rm max}$. This is computationally more expensive but yields the advantage that we can estimate the error due to the restriction of the bond dimension by adding up the weight from discarded singular values.
The MPS representation of the state gives cheap access to all Rényi entropies, which are calculated from the singular spectra of the corresponding cuts. The second cumulant and covariance can be calculated from expectation values of bond dimension two- and three MPOs respectively. The computational effort for calculating these observables is negligible compared to the unitary time evolution which costs $O(Lk\chi^{2}t_{\rm max}/\text{min}\{\delta t,t_{\gamma}\})$ [134, 135] where $t_{\gamma}=1/(L\gamma)$ is the average time between two measurements.

### v2

@@ VII.1 Summary of main results
In the absence of interaction (and for measurements preserving Gaussianity of the theory), this formalism allows one to identify the symmetry class of the problem in terms of the symmetry of the Lagrangian, Sec. II.6. For complex fermions, this places the problem in the symmetry class AIII or BDI, depending on the presence of additional particle-hole symmetry of Lagrangian, in agreement with earlier results. We have derived the replicated Keldysh non-linear sigma model for both symmetry classes, Sec. III.1. The model incorporates boundary conditions (Fig. 2) specifically designed to produce generating functions for charge cumulants and Rényi entropies directly in the language of the sigma model (Sec. III.2).
@@
We have corroborated these analytical results with a numerical study of an interacting 1D model (Sec. VI). Since exact diagonalization proved insufficient in view of system-size limitations, we have employed the time-dependent variational principle for matrix product states, Sec. VI.1) enabling us to investigate significantly larger systems. The numerically observed behavior of both the entanglement entropy and charge cumulants (Figs. 10 and 11) aligns very well with our analytical predictions. In particular, our MPS-TDVP numerics demonstrate the separation between information and charge, characteristic of the phase with “delocalized” information and charge fluctuations. Furthermore, the behavior of the corresponding behavior of entanglement and charge cumulant is consistent with the predicted area-law and logarithmic-law scaling, respectively. Our estimated values for critical $\gamma$ for both transitions show that two transitions are located quite close to each other. In addition, we have conducted time-dependent Hartree-Fock simulations (Sec. VI.2 and Fig. 12) as a complementary method to estimate the transition point between the area-law phase and the logarithmic phase for charge fluctuations. While the numerical data suggest that the entanglement transition takes place at a somewhat larger value of $\gamma$ than the charge-fluctuation transition (as in the phase diagram in Fig. 9), the accuracy is not sufficient to deduce this with certainty.
@@ VII.2 Comparison with related works
In a number of works [88, 89, 90, 91, 93, 94, 95], the entanglement in 1D models of interacting bosons subjected to generalized measurements was studied numerically. These works reported evidence of a transition from area-law to volume-law scaling of the entanglement entropy. While some of these models were different from our interacting-fermion model, and the measurement protocols were also different, one can expect that the physics that we address has a large degree of universality. In this sense, the conclusions of the above works on the area-law-to-volume-law entanglement transition are in agreement with our results. However, the scaling of particle-number fluctuations for interacting bosons, as well as the associated charge-sharpening transition, had not been explicitly addressed in these past papers.
@@ VII.3 Outlook
    The approach can also be directly generalized to describe a monitored open system in contact with environment, whose evolution includes measurements and non-unitary Lindbladian dynamics (see, e.g., Refs. [141, 142] for review). For such systems, however, the entanglement entropy and particle number cumulant cannot serve as good observables to define the transition as their behavior is expected to follow generically a volume law due to effect of the environment. On the other hand, the volume-law contribution of “classical correlations” is expected to cancel [143] in the scaling of the mutual information and particle-number covariance, making them suitable observables to study possible MIPTs.
  • (iii)
@@
    Another important direction is to extend the theory to study dynamics (i.e., correlation functions with different time arguments. This will allow one, in particular, to directly explore the dynamics of purification and charge sharpening.
  • (v)
@@ Acknowledgements.
We are very grateful to Elmer Doggen, Adam Nahum, Pavel Ostrovsky, Marcin Szyniszewski, and Romain Vasseur for insightful discussions. Note added: While preparing the manuscript for publication, we became aware of related work [144] on monitored interacting fermions. We thank Matthew Foster for informing us about this work.
@@ Appendix C Numerical comparison of weak-localization corrections in AIII and BDI classes
In Fig. 13, we plot the behavior of the running coupling constant for both symmetry classes (left panel: AIII, right panel: BDI), as obtained from the density correlation function. The dependence of $g(q)$ at large momenta corresponds to the ballistic behavior: $C(q)\propto\operatorname{const}\Rightarrow g(q)\propto 1/q$; the dashed line marks a universal interpolating function between ballistic ($q\ell_{0}\gtrsim 1$) and diffusion ($q\ell_{0}\ll 1$) regions. Evident bending down of each curve is a manifestation of localization. It is evident that the qualitative behavior of the curves in the two panels of Fig. 13 is identical. At the same, a closer inspection reveals that the localization effects are stronger in the BDI model than in the AIII model, in consistency with analytical predictions
@@ Appendix E Calculation of the Hartree and Fock diagrams for the interaction Lagrangian
As a consequence, at $z\to 0$, the $\tau$-integral in Eq. (217) logarithmically diverges for $d=1$ and saturates at a constant of order unity for $d>1$, yielding Eq. (104) of the main text.
The behavior of $Y_{\text{HF}}$ for $d=1,2,3$ is plotted in Fig. 16.
@@ I.1 Time-dependent variational principle
For each bond expansion step, we allow for $k=0.3\chi$ extra singular vectors, expanding a bond of dimension $\chi$. This is somewhat larger than $k=0.1\chi$ discussed in Ref. [136] in the context of density matrix renormalization group. We found the additional SVD necessary and kept the suggested oversampling parameter of $p\leq 10$. If a maximum bond dimension $\chi_{\rm max}$ is imposed, we still allow for an expansion of any bond beyond $\chi_{\rm max}$ within these expansion parameters at every time step. If the dimension of the expanded bond exceeds $\chi_{\rm max}$ it is truncated back to $\chi_{\rm max}$. This is computationally more expensive but yields the advantage that we can estimate the error due to the restriction of the bond dimension by adding up the weight from discarded singular values.
The convergence with respect to the bond dimension is shown in Fig. 17 where the entropy $S_{\ell}$ is presented for our largest system size $L=72$, strongest interaction $V=1$, and the weakest measurement rate $\gamma=0.3$, i.e., for the “worst case” from the point of view of MPS simulations.
The MPS representation of the state gives cheap access to all Rényi entropies, which are calculated from the singular spectra of the corresponding cuts. The second cumulant and covariance can be calculated from expectation values of bond dimension two- and three matrix product operators respectively. The computational effort for calculating these observables is negligible compared to the unitary time evolution which costs $O(Lk\chi^{2}t_{\rm max}/\text{min}\{\delta t,t_{\gamma}\})$ [135, 136] where $t_{\gamma}=1/(L\gamma)$ is the average time between two measurements.
@@
Figure 17: Entanglement entropy $S_{\ell}$ in a system of size $L=72$, as a function of cut position $\ell$, for bond dimensions $\chi_{\text{max}}=128,256$, and 512, the interaction strength $V=1$ and the measurement rate $\gamma=0.3$.